OpenStudy (anonymous):

A football quarterback has 2 more chances to throw a touchdown before his team is forced to punt the ball. He misses the receiver on the first throw 30% of the time. When his first throw is incomplete, he misses the receiver on the second throw 10% of the time. What is the probability of not throwing the ball to a receiver on either throw?

OpenStudy (anonymous):

3/20 would be the right answer I think, but i don't know for sure.

OpenStudy (anonymous):

or 15%

OpenStudy (kropot72):

Let A be the event he misses on the first throw. P(A) = 0.3 Let B be the event he misses on the second throw. Then P(B|A) = 0.1 \[\large P(A\cap B)=P(A)\times P(B|A)=0.3\times0.1=0.03\]

OpenStudy (kropot72):

If the probability of not throwing the ball to a receiver on either throw is expressed as a percentage it is 0.03 * 100 = 3%.

OpenStudy (anonymous):

are you sure that is right how did you get that? @Pencera

OpenStudy (anonymous):

no im not i said that in my answer

OpenStudy (anonymous):

i apologize